Equipment was purchased for $140000.

Kimmel, Accounting, 6e S OF ACCOUNTING (ACCT 2101, 2102, Assignment Gradebook ORION Downloadable eTextbook Question 15 Equipment was purchased for $1405oo. Freight charges amounted to $7000 and there was a cost of $17000 for building a foundation and installing the equipment. It s estimated that the equipment will have a $34000 salvage value at the end of its S-year useful life. Depreciation Expense each year using the straight-line method will be D $32800. D $21200 D $23200o : с $26000. Open Show Work Click if you would like to Show Work for this question:
Answer #1

Total Cost of equipment = 140000+7000+17000 = 164000

Straight line depreciation = (164000-34000/5) = 26000

So answer is d) $26000